# Biases in gendered citation practices: an exploratory study and some reflections on the Matthew and Matilda effects

### v1

@@ Abstract
The number of citations of scientific articles has a huge impact on recommendations for funding allocations, recruitment decisions, promotion decisions and awards, just to name a few. Recent studies conducted in different scientific disciplines (e.g., physics and neuroscience) have concluded that researchers belonging to some socio-cultural groups (e.g., women, racialized people) are usually less cited than other researchers belonging to dominating groups. This is usually due to the presence of citation biases in reference lists. These citation biases towards researchers from some socio-cultural groups may inevitably cause unfairness and inaccuracy in the assessment of articles impact. These citation biases may therefore translate to significant disparities in promotion, retention, grant funding, awards, collaborative opportunities, and publications. In this paper, we conduct –to the best of our knowledge – the first study aiming at analyzing gendered citation practices in the software engineering (SE) literature. Our study allows reflecting on citations practices adopted in the SE field and serves as a starting point for more robust empirical studies on the analyzed topic. Our results show that some efforts still need to be done to achieve fairness in citation practices in the SE field. Such efforts may notably consist in the inclusion of citation diversity statements in manuscripts submitted for publication in SE journals and conferences. Such efforts may also consist in redefining the power dynamics across scientific communities, industry, and academia to foster scientific inclusion (e.g., fair representation of scientific contributions) in the SE field.
@@ I Introduction
A citation bias occurs when the authors of a given paper decide to include or not to include in their reference list a given reference based on considerations that are not related to the reference relevance and its quality [25]. Citation biases are detrimental to various socio-cultural groups, especially women and racialized researchers since their papers are usually under-cited [3, 12]. For instance, in the most prolific countries, the papers in which women are the lead authors are usually less cited than the ones that have men as lead authors [1]. Thus, citation biases aggravates gender disparities especially since citations are key in the assessment of researchers [1]. That gender disparity is very pervasive and translates in academia in terms of research performance, the recruitment of faculty members, the number of citations researchers receive, grant applications, determination of the positions of authors in a paper, and the choice of academic awardees, just to name a few [2]. Besides, male researchers are more advantaged than female researchers when it comes to recruitment and promotion [2]. Furthermore, male researchers are usually more credited than female researchers when it comes to the contributions they make to research [2, 8]. This may give the false impression that female researchers are less capable than male researchers. This could notably be detrimental to the academic career of women [2] and to the innovation, creativity, performance, and productivity in various scientific and/or technological projects [18, 29, 30, 36, 39, 42]. That innovation is particularly key in the software engineering (SE) field since it drives technological advances, which influences various aspects of the society as a whole. Such aspects include the economy, the environment, as well as politics, just to name a few.
The disparity of female in STEM (science, technology, engineering, and mathematics) disciplines is well-established [1, 3, 30, 31, 35, 36, 38, 40, 41, 46]. That disparity is well-pronounced in the software industry and has been increasingly explored (e.g., [35, 38, 40, 41, 42, 48, 49, 50]). Still, when it comes to research in particular, the gap in citations that researchers belonging to some socio-cultural groups may experience and that may stem from citation biases remains understudied [3]. Besides, the awareness about that topic remains low [3]. This makes it difficult to devise novel solutions to mitigate citation biases and make sure the scientific papers we publish sufficiently reflect the intellectual diversity of the existing scientific contributions.
@@
  • Contribution 3: we provide key insights on the fairness of citations practices in the SE field.
@@ II-B Citation Biases in Literature
Gender: Wu [44] surveyed research focusing on the links between gender and citations. Several studies (e.g., [1, 6, 9, 11, 14, 16, 19, 21, 39, 43, 45, 47]) consistently highlighted persistent gender citation biases. These studies usually concluded that, despite an increase in women authorship, women-authored papers continue to receive fewer citations than men-authored papers. For instance, in economics, Ferber and Brun [16] found that male-only papers receive more citations than female-only or mixed-gender authored papers. Teich et al. [14] reported similar trends in physics, with male-authored papers being over-cited and women-authored papers under-cited, varying by sub field and proximity[14]. Larivière et al. [1] found that women in prominent author positions (sole, first and last) attracted fewer citations and consistently had more domestic portfolios with fewer international collaborations than men. Dworkin et al. [11] showed that although women’s representation increased in neuroscience, citation disparities persisted. Fulvio et al. [9] found that male-authored papers were still over-cited in cognitive neuroscience despite an increase in women-authored papers.
@@ II-C Citation Bias Detection Methods and Their Limitations
    Collect metadata on papers from top journals; this approach is popular in research which focuses on a single field [11, 5, 16, 14].
@@
    Collect metadata on a large set of papers from a variety of journals; this is an approach used for research interested in global citation patterns [17, 1].
@@
Data Pre-processing: This usually allows labelling authors and authors cited within each paper based on some interest criteria i.e attributes. Studies on gender bias used databases such as the US Census, Wikipedia, and Social Security Administration to match names to gender [16, 14, 1, 11, 5, 9]. Bertolero et al. [5] assigned race using probabilistic databases and neural networks trained on American Voter and Census data. Geographic bias studies such as Gomez et al. [17] extracted author location information. Duyx et al. [15] extracted information from the papers content, which included information such as the number of positive articles, the number of negative articles, the number of citations to positive articles, and the number of citations to negative articles. Noteworthy, at this stage, the removal of self-citations from reference lists is common practice [11, 5, 16, 14, 9, 22].
@@
Limitations: Studies which extract data from top journals must be careful with generalizing bias within the field as a whole [11, 5, 14]. Dworkin et al. [11] state that they did not account for authors’ institutional prestige, potentially introducing bias due to gender imbalance in hiring and prestige-based citation behaviour. Gender-based papers such as Dworkin et al. [11] caution that gender determination methods are binary, excluding intersex, transgender, and non-binary identities. Bertolero et al. [5] investigated race bias, which is limited by probabilistic racial and ethnic identity analyses based on limited racial categories and outdated census data that may inaccurately assign categories. Furthermore, Gomez et al. [17] acknowledge that their citational lensing framework is limited in accuracy as the textual similarity between countries is inherently noisy, which impacts comparisons with citation data. Duyx et al. [15]’s meta-study on result-biased citation acknowledges the need for caution in making broad conclusions about citation bias across all sciences due to the high heterogeneity in their analyses.
@@ II-D Citation Biases Mitigation Methods and their Limitations
The literature proposed several solutions to mitigate citation biases. For instance, Larivière et al. [1, 26] detected global gender bias against women and recommended fostering international collaboration programs for women researchers to reduce disparities in research output and impact. Additionally, they state that policymakers should consider the diverse contexts in which science is performed and identify mechanisms that perpetuate gender inequality.
Teich et al. [14] found that women are cited more in longer reference lists, thus noting that removing length limits on reference lists could promote gender equality. However, this should not imply that man-authored papers are more valuable. Bertolero et al.[5] recognized racial citation bias against authors of colour, and, along with Dworkin et al. [11] they advocate for education, transparency, personal responsibility, and ally-ship within research communities. Additionally, they emphasize that non-minority male scholars must recognize and address the challenges they create and perpetuate.
Duyx et al. [15] stated that journals should include declarations about the representativeness of the cited literature, akin to those for funding and author contributions. This could help raise awareness on selective citation practices and mitigate result-based citation bias. Along the same lines, several authors such as Rowson et al. [10], Dworkin et al. [11], Teich et al. [14], and Ray et al. [25] advocate for the inclusion Citation Diversity Statements (CDSs) in papers. A CDS is a paragraph added before the reference list of a paper [25]. It shows a commitment to citation equity [25]. Full CDSs provide a percentage breakdown of citations, explain assessment methods and limitations, and highlight citation diversity [10, 25].
@@ III-B Step 1: Data Collection
We stored the Plain Text Files in folders by journal and imported into Dworkin’s R codebase for processing. We stored the post processed dataset in a MongoDB which we use later for the statistical analysis. Our raw data and processed dataset are available online  [^3] Link to the Raw data, see /R-processing/wos-data  [^4] Link to the MongoDB dataset exported as a set of JSON files.
@@ III-C Step 2: Name Processing
We use Dworkin et al. [11]’s R code to process the metadata we collected in Step 1. In their code, Dworkin et al. [11] cross-reference any initial only names as provided by WoS with CrossRef by querying the CrossRef API with the article DOI. If the requests return a full name, then the WoS name is replaced. Secondly, Dworkin et al. implemented a name disambiguation algorithm which matches name variants to connect papers authored by the same person, but are published under different names. This algorithms functions by matching all entries which share the first name initial and/or middle initial and last name are the same. Since, last names are matched first, followed by matching first and/or middle initials, it is crucial to ensure all last names are complete. In Dworkin et al. [11]’s words “For example, if an entry listed an author as R. J. Dolan, and we found matches under Ray J. Dolan and Raymond J. Dolan, we would replace the R. J. Dolan entry with the more common completed variant. If, instead, we found matches under Ray J. Dolan and Rebecca J. Dolan, we would not assign a name to the original R. J. Dolan entry.” (pg. 11).
Dworkin et al. [11] noted in their database all last names where complete in the initial WoS metadata. However, these was not the case for our dataset, therefore, we implemented an additional R script which removed any articles which has either initial or missing last names. In total, we removed 853 articles which had incomplete last names. Thus, this reduced our dataset size to 160,808 articles. The Secure Open Enterprise Master Patient Index is a scientific record linkage tool which contains a databases of common nicknames. We used it to connect name variants [21]. Any names where the full names cannot be matched or retrieved with CrossRef will not be assigned a gender and is excluded from the analysis. In this step, CrossRef replaced a total of 19,164 names. Besides, 14,474 name variations matches were found.
@@ III-D Step 3: Gender Assignment
  • First round of gender assignment: in this round, the Social Security Administration (SSA) database as implemented in the gender R package is used to assign genders to authors whose first names are available in the processed dataset. This SSA database provides the proportion of male-female infants assigned to a name in the US from 1932 to 2012. If a given author’s name yields a proportion greater or equal to 0.7, then a gender is assigned to the corresponding author. Noteworthy, picking 0.7 as a threshold is common the literature (e.g., [9, 11, 43]).
@@ III-E Step 4: Citation Processing
Following Dworkin et al. [11]’s methodology, for our analysis we do not consider all articles cited in the metadata reference lists, instead we only consider citations of articles which are in our dataset. Thus, we must identify any articles present in the metadata references lists that also exists in our dataset; all other referenced articles will be removed from the reference list. We explain this process below.
@@
As we mention in Section II, the removal of self-citations from reference lists is a well-established practice in citation analysis (e.g., [11, 5, 16, 14, 9, 22]). It allows mitigating the impact gender differences may have on self-citation patterns [9]. Dworkin et al. [11] define a self-citation as any paper listed in a reference list where the first or last author of that paper is also the first or last author of the citing paper. Dworkin et al. acknowledge that this is a restrictive definition. However, due to the nature of this analysis, since cited genders will be based on assignments already present in our database, this is only way we can guarantee the authors are present in our database. We rely on Dworkin et al. [11]’s code to identify potential self-citations. This process consists of collecting a list of all first authors and a list of all last authors. These names are then matched to any papers where the authors name appear in that list. Each paper stores a list of indices which point to the papers in our database which have the same author. These indices represent potential self-citations.
@@
We export the processed R Data to a MongoDB database. The processed dataset has now been gender labelled and indexed with pointers that connect cited papers, to the citing paper as well as papers authored by the same person. This section of the Methodology workflow is where we deviate away from Dworkin et al. [11]’s statistical analysis that leverages a Generalized Additive Model (GAM), and adapt a simpler analysis similar to Fulvio et al. [9]’s. This allows ensuring the citation processing step is suitable for the journals we target. Hence, in accordance with Fulvio et al. [9], we rely on our own Python script to automatically remove the self-citations we identified using the code of Dworkin et al. [11]. Thus, we consider that each article in the processed dataset includes a list of cited article indices and a list of article indices which are authored by the same person. If an article index appears in both lists, our script removes it from the cited list as it means it is a self citation. Besides, when assigning gender categories to cited articles, we use the list of cited article indices for each article in our database to create a new list which stores gender categories instead of indices.
@@
Since, the authors which were not able to be assigned gender were simply labeled as Unknown but not immediately removed from the dataset in Step 3, the citation processing was conducted on the processed dataset (160,808 articles). In total 567,374 citations were retrieved and 38,243 of those were removed as they were self-citations. However, these citations include citing and cited articles which contain authors without assigned genders. Thus, when we filter those out of the dataset, we obtained 107,152 articles and 280,217 citations of which 25,561 were removed as self-citations. The citation data was very sparse up until 2009, with only 3,482 citations in total for those 5 years (1.37 % of all citations). Thus, in accordance with the literature (e.g., [6, 9, 11]), we decided at this stage to conduct our citation analysis for articles published from January 2009- August 2024. This will allow us in Step 5 to focus on citations that yield a more robust statistical citation analysis. In the current Step (i.e. Step 4), when filtering out articles published from 2003-2008, we ended up with 96,282 articles and 276,187 citations of which 25,013 are self-citations. Thus, our final citation analysis dataset consists of 96,282 articles with 251,174 citations.

### v3

@@ Abstract
The number of citations of scientific articles has a huge impact on recommendations for funding allocations, recruitment decisions, promotion decisions and awards, just to name a few. Recent studies conducted in different scientific disciplines (e.g., physics and neuroscience) have concluded that researchers belonging to some socio-cultural groups (e.g., women, racialized people) are usually less cited than other researchers belonging to dominating groups. This is usually due to the presence of citation biases in reference lists. These citation biases towards researchers from some socio-cultural groups may inevitably cause unfairness and inaccuracy in the assessment of articles impact. These citation biases may therefore translate to significant disparities in salaries, promotion, retention, grant funding, awards, collaborative opportunities, and publications. In this paper, we conduct –to the best of our knowledge – the first study aiming at analyzing gendered citation practices in the software engineering (SE) literature. Our study allows reflecting on citation practices adopted in the SE field and serves as a starting point for more robust empirical studies on the analyzed topic. Our results show that some efforts still need to be done to achieve fairness in citation practices in the SE field. Such efforts may notably consist in the inclusion of citation diversity statements in manuscripts submitted for publication in SE journals and conferences. Such efforts may also consist in redefining the power dynamics across scientific communities, industry, and academia to foster scientific inclusion (e.g., fair representation of scientific contributions) in the SE field.
@@ I Introduction
A citation bias occurs when the authors of a given paper decide to include or not to include in their reference list a given reference based on considerations that are not related to the reference relevance and its quality [25]. Citation biases are detrimental to various socio-cultural groups, especially women and racialized researchers since their papers are usually under-cited [3, 12, 54]. For instance, in the most prolific countries, the papers in which women are the lead authors are usually less cited than the ones that have men as lead authors [1]. Thus, citation biases may aggravate gender disparities especially since citations are key in the assessment of researchers [1]. That gender disparity is very pervasive and translates in academia in terms of research performance, the recruitment of faculty members, the number of citations researchers receive, grant applications, determination of the positions of authors in a paper, and the choice of academic awardees, just to name a few [2]. Besides, male researchers are more advantaged than female researchers when it comes to recruitment and promotion [2]. Furthermore, male researchers are usually more credited than female researchers when it comes to the contributions they make to research [2, 8]. This may give the false impression that female researchers are less capable than male researchers. This could notably be detrimental to the academic career of women [2] and to innovation, creativity, performance, and productivity in various scientific and/or technological projects [18, 29, 30, 36, 39, 42]. That innovation is particularly key in the software engineering (SE) field since it drives technological advances, which influences various aspects of the society as a whole. Such aspects include the economy, the environment, as well as politics, just to name a few.
The disparity of women in STEM (science, technology, engineering, and mathematics) disciplines is well-established [1, 3, 30, 31, 36, 38, 40, 41, 46, 55]. That disparity is well-pronounced in the software industry and has been increasingly explored (e.g., [35, 38, 40, 41, 42, 48, 49, 50]). Still, when it comes to research in particular, the gap in citations that researchers belonging to some socio-cultural groups may experience and that may stem from citation biases remains understudied [3]. Furthermore, the awareness about that topic remains low [3]. This makes it difficult to devise novel solutions to mitigate citation biases and make sure the scientific papers we publish sufficiently reflect the intellectual diversity of the existing scientific contributions.
@@
  • Contribution 3: we provide key insights on the fairness of citation practices in the SE field.
@@ II-B Citation Biases in Literature
Gender: Wu [44] surveyed research focusing on the links between gender and citations. Several studies (e.g., [1, 6, 9, 11, 14, 16, 19, 21, 39, 43, 45, 47]) highlighted persistent gender citation biases. These studies usually concluded that, despite an increase in women authorship, women-authored papers continue to receive fewer citations than men-authored papers. For instance, in economics, Ferber and Brun [16] found that male-only papers receive more citations than female-only or mixed-gender authored papers. Teich et al. [14] reported similar trends in physics, with male-authored papers being over-cited and women-authored papers under-cited, varying by sub field and proximity[14]. Larivière et al. [1] found that women in prominent author positions (sole, first and last) attracted fewer citations and consistently had more domestic portfolios with fewer international collaborations than men. Dworkin et al. [11] showed that although women’s representation increased in neuroscience, citation disparities persisted. Fulvio et al. [9] found that male-authored papers were still over-cited in cognitive neuroscience despite an increase in women-authored papers.
@@ II-C Citation Bias Detection Methods and Their Limitations
    Collect metadata on papers from top journals; this approach is popular in research which focuses on a single field (e.g., [11, 5, 16, 14]).
@@
    Collect metadata on a large set of papers from a variety of journals; this is an approach used for research interested in global citation patterns (e.g., [17, 1]).
@@
Data Pre-processing: This usually allows labelling authors and authors cited within each paper based on some interest criteria i.e attributes. Studies on gender bias used databases such as the US Census, Wikipedia, and Social Security Administration to match names to gender [16, 14, 1, 11, 5, 9]. Bertolero et al. [5] assigned race using probabilistic databases and neural networks trained on American Voter and Census data. Geographic bias studies such as the one of Gomez et al. [17] extracted author location information. Duyx et al. [15] extracted information from the papers content, which included information such as the number of positive articles, the number of negative articles, the number of citations to positive articles, and the number of citations to negative articles. Note that, at this stage, the removal of self-citations from reference lists is common practice [11, 5, 16, 14, 9, 22].
@@
Limitations: Studies which extract data from top journals must be careful with generalizing bias within the field as a whole [11, 5, 14]. Dworkin et al. [11] state that they did not account for authors’ institutional prestige, potentially introducing bias due to gender imbalance in hiring and prestige-based citation behaviour. Gender-based papers such as Dworkin et al. [11] caution that gender determination methods are binary, excluding intersex, transgender, and non-binary identities. Bertolero et al. [5] investigated race bias, which is limited by probabilistic racial and ethnic identity analyses based on limited racial categories and outdated census data that may inaccurately assign categories. Furthermore, Gomez et al. [17] acknowledge that their citational lensing framework is limited in accuracy as the textual similarity between countries is inherently noisy, which impacts comparisons with citation data. Duyx et al. [15]’s meta-study on result-biased citations acknowledges the need for caution in making broad conclusions about citation bias across all sciences due to the high heterogeneity in their analyses.
@@ II-D Citation Biases Mitigation Methods and their Limitations
The literature proposed several solutions to mitigate citation biases. For instance, Larivière et al. [1, 26] detected global gender bias against women and recommended fostering international collaborations programs for women researchers to reduce disparities in research output and impact. Additionally, they state that policymakers should consider the diverse contexts in which science is performed and identify mechanisms that perpetuate gender inequality.
Teich et al. [14] found that women are cited more in longer reference lists, thus noting that removing length limits on reference lists could promote gender equality. However, this should not imply that man-authored papers are more valuable. Bertolero et al.[5] recognized racial citation bias against authors of colour, and, along with Dworkin et al. [11], they advocate for education, transparency, personal responsibility, and ally-ship within research communities. Additionally, they emphasize that non-minority male scholars must recognize and address the challenges they create and perpetuate.
Duyx et al. [15] stated that journals should include declarations about the representativeness of the cited literature, akin to those for funding and author contributions. This could help raise awareness on selective citation practices and mitigate result-based citation bias. Along the same lines, several authors such as Rowson et al. [10], Dworkin et al. [11], Teich et al. [14], and Ray et al. [25] advocate for the inclusion of Citation Diversity Statements (CDSs) in papers. A CDS is a paragraph added before the reference list of a paper [25]. It shows a commitment to citation equity [25]. Full CDSs provide a percentage breakdown of citations, explain assessment methods and limitations, and highlight citation diversity [10, 25].
@@ III-B Step 1: Data Collection
We stored the Plain Text Files in folders by journal and imported into Dworkin’s R codebase for processing. We stored the post processed dataset in MongoDB –a document-oriented database – which we use later for the statistical analysis. Our raw data and processed dataset are available online  [^3] Link to the Raw data, see /R-processing/wos-data  [^4] Link to the MongoDB dataset exported as a set of JSON files.
@@ III-C Step 2: Name Processing
We use Dworkin et al. [11]’s R code to process the metadata we collected in Step 1. In their code, Dworkin et al. [11] cross-reference any initial only names as provided by WoS with CrossRef by querying the CrossRef API with the article DOI. If the requests return a full name, then the WoS name is replaced. Secondly, Dworkin et al. implemented a name disambiguation algorithm which matches name variants to connect papers authored by the same person, but are published under different names. This algorithm functions by matching all entries which share the first name initial and/or middle initial and last name are the same. Since, last names are matched first, followed by matching first and/or middle initials, it is crucial to ensure all last names are complete. In Dworkin et al. [11]’s words “For example, if an entry listed an author as R. J. Dolan, and we found matches under Ray J. Dolan and Raymond J. Dolan, we would replace the R. J. Dolan entry with the more common completed variant. If, instead, we found matches under Ray J. Dolan and Rebecca J. Dolan, we would not assign a name to the original R. J. Dolan entry.” (pg. 11).
Dworkin et al. [11] noted in their database all last names where complete in the initial WoS metadata. However, these was not the case for our dataset, therefore, we implemented an additional R script which removed any articles which has either initial or missing last names. In total, we removed 853 articles which had incomplete last names. Thus, this reduced our dataset size to 160,808 articles. The Secure Open Enterprise Master Patient Index is a scientific record linkage tool which contains a database of common nicknames. We used it to connect name variants [21]. Any names where the full names cannot be matched or retrieved with CrossRef will not be assigned a gender and is excluded from the analysis. In this step, CrossRef replaced a total of 19,164 names. Besides, 14,474 name variations matches were found.
@@ III-D Step 3: Gender Assignment
  • First round of gender assignment: in this round, the Social Security Administration (SSA) database as implemented in the gender R package is used to assign genders to authors whose first names are available in the processed dataset. This SSA database provides the proportion of male-female infants assigned to a name in the US from 1932 to 2012. If a given author’s name yields a proportion greater or equal to 0.7, then a gender is assigned to the corresponding author. Noteworthy, picking 0.7 as a threshold is common in the literature (e.g., [9, 11, 43]).
@@ III-E Step 4: Citation Processing
Following Dworkin et al. [11]’s methodology for our analysis, we do not consider all articles cited in the metadata reference lists, instead we only consider citations of articles which are in our dataset. Thus, we must identify any articles present in the metadata references lists that also exists in our dataset; all other referenced articles will be removed from the reference list. We explain this process below.
@@
As we mention in Section II, the removal of self-citations from reference lists is a well-established practice in citation analysis (e.g., [11, 5, 16, 14, 9, 22]). It allows mitigating the impact gender differences may have on self-citation patterns [9]. Dworkin et al. [11] define a self-citation as any paper listed in a reference list where the first or last author of that paper is also the first or last author of the citing paper. Dworkin et al. acknowledge that this is a restrictive definition. However, due to the nature of this analysis, since cited genders will be based on assignments already present in our database, this is the only way we can guarantee the authors are present in our database. We rely on Dworkin et al. [11]’s code to identify potential self-citations. This process consists of collecting a list of all first authors and a list of all last authors. These names are then matched to any papers where the authors name appear in that list. Each paper stores a list of indices which point to the papers in our database which have the same author. These indices represent potential self-citations.
@@
We export the processed R Data to a MongoDB database. The processed dataset has now been gender labelled and indexed with pointers that connect cited papers, to the citing paper as well as papers authored by the same person. This section of the Methodology workflow is where we deviate away from Dworkin et al. [11]’s statistical analysis that leverages a Generalized Additive Model (GAM), and adapt an analysis similar to Fulvio et al. [9]’s. This allows ensuring the citation processing step is suitable for the journals we target. Hence, in accordance with Fulvio et al. [9], we rely on our own Python script to automatically remove the self-citations we identified using the code of Dworkin et al. [11]. Thus, we consider that each article in the processed dataset includes a list of cited article indices and a list of article indices which are authored by the same person. If an article index appears in both lists, our script removes it from the cited list as it means it is a self citation. Besides, when assigning gender categories to cited articles, we use the list of cited article indices for each article in our database to create a new list which stores gender categories instead of indices.
@@
Since, the authors which were not able to be assigned gender were simply labeled as Unknown but not immediately removed from the dataset in Step 3, the citation processing was conducted on the processed dataset (160,808 articles). In total 567,374 citations were retrieved and 38,243 of those were removed as they were self-citations. However, these citations include citing and cited articles which contain authors without assigned genders. Thus, when we filter those out of the dataset, we obtained 107,152 articles and 280,217 citations of which 25,561 were removed as self-citations. The citation data was very sparse up until 2009, with only 3,482 citations in total for those five years (i.e. 1.37 % of all citations). Thus, in accordance with the literature (e.g., [6, 9, 11]), we decided at this stage to conduct our citation analysis for articles published from January 2009- August 2024. This will allow us in Step 5 to focus on citations that yield a more robust statistical citation analysis. In the current Step (i.e. Step 4), when filtering out articles published from 2003-2008, we ended up with 96,282 articles and 276,187 citations of which 25,013 are self-citations. Thus, our final citation analysis dataset consists of 96,282 articles with 251,174 citations.